OpenStudy (anonymous):

Drawing an embarrassing blank. How exactly do I solve this. (6x-1)^2<1

OpenStudy (anonymous):

square, subtract 1, solve a quadratic

OpenStudy (anonymous):

or you could start with \[-1\leq 6x-1\leq 1\] which might be easier

OpenStudy (anonymous):

since \[a^2<1\iff -1<a<1\]

OpenStudy (anonymous):

I some how tried that but didnt get it right

OpenStudy (anonymous):

\[-1< 6x-1<1\] \[0<6x<2\] \[0<x<\frac{1}{3}\] should work, is that not it?

OpenStudy (anonymous):

ah thank god I found out what I was doing wrong

OpenStudy (anonymous):

something silly no doubt

OpenStudy (anonymous):

you are a hero.. so used to complex math forget the basics